OpenStudy (osanseviero):

I need help (and explanation) of how to solve this expresion, giving an answer without negative exponents

OpenStudy (anonymous):

SUre. What's the problem?

OpenStudy (osanseviero):

\[\left( \left( -4x ^{4}y ^{-2} \right)\div \left( 5x ^{-1}y ^{4} \right) \right)^{-4}\] You dont need to solve

OpenStudy (osanseviero):

Just explain me how to do it (which is the order)

OpenStudy (anonymous):

Basically, whenever you have a negative exponent when there's one term over another, then the terms switch places from top to bottom.

OpenStudy (anonymous):

First distribute the -4 exponent.

OpenStudy (osanseviero):

ok...shouldnt it be first to divide?

OpenStudy (anonymous):

I guess you could. Yeah. It would be easier that way. Start inside out.

OpenStudy (anonymous):

Just set up a huge fraction to start off.

OpenStudy (precal):